OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 15 degrees Celsius is. PLEASE EXPLAIN HOW TO DO IT

OpenStudy (anonymous):

\[C=\frac{5}{9}(F-32)\]to solve for F, you have to move 5/9 and 32 to the other side, do you know how to do it?

OpenStudy (anonymous):

No haha Its kinda Confusing :P lol

OpenStudy (ganpat):

step 1. F = (9/5) C + 32... equation for F & c.. Step 2.. now substitute C = 15 in given equation.. to find F..

OpenStudy (ganpat):

F = (9/5) C + 32 F = (9/5) (15) + 32 F = 9 * 3 + 32 F = 27 + 32 F = 55 F....

OpenStudy (anonymous):

Could You explain each step in detail please? I still dont get it sorry :P

OpenStudy (ganpat):

ok.. Step 1... is the standard formula.. which we cannot edit much... it relates between two different units... i.e. Fahrenheit and degree.. Degree is International standard unit and Fahrenheit is US standard units..

OpenStudy (anonymous):

Okay:)

OpenStudy (ganpat):

while in step 2.. i have substituted value of C i.e. in the standard equation.. and calculated for F.. its just simplification...

OpenStudy (anonymous):

Okay

OpenStudy (ganpat):

understood ???

OpenStudy (anonymous):

Okay im with you:)

OpenStudy (ganpat):

its done.. :-p..

OpenStudy (anonymous):

F = (9/5) C + 32 F = (9/5) (15) + 32 F = 9 * 3 + 32 F = 27 + 32 F = 55 F.... When you did that how did you get 9*3 + 32??

OpenStudy (ganpat):

F = (9/5) C + 32 ..... standard formula.. F = (9/5) (15) + 32 ..... substitute C = 15, from the question.. F = 9 * 3 + 32 ..... 15 is divided by 5... so the answer is 3 F = 27 + 32 ....... 9 * 3 = 27.... F = 55 F ...... 27 and 32 are added..

OpenStudy (ganpat):

@Ladiddaaaa : still confusion ??

OpenStudy (anonymous):

ohhhhhhh k thnx :D :D :D :D :D

OpenStudy (ganpat):

w/c...